OpenStudy (ja1):

Brian is creating a collage on a piece of cardboard that has an area of 150r^2 square centimeters. The collage is covered entirely by pieces of paper that do not overlap. Each piece has an area of √r^3 square centimeters. Use the given information to determine an expression for the total number of pieces of paper used.

OpenStudy (ja1):

\[\frac{ 150r^2 }{ \sqrt{r^3} }\] I made this^^

OpenStudy (ja1):

That's √r^3 btw

OpenStudy (ja1):

@mathstudent55 @timo86m

OpenStudy (ja1):

@thomaster @dan815

OpenStudy (luigi0210):

I saw a problem somewhat similar to this earlier..

OpenStudy (ja1):

Really? Maybe same quiz xD anyways know how to do this?

OpenStudy (ja1):

my last question and I wanna' sleep D:

OpenStudy (luigi0210):

http://openstudy.com/users/unheard#/updates/52100ccde4b0450ed75cf185 This is it, might help, might not, but oh well worth a try.

OpenStudy (ja1):

ugh all that rationalizing stuff, I was meaning to ask how to do it but I can't bother to learn it now xD do you know any other methods for this?

mathslover (mathslover):

You don't know rationalizing?

OpenStudy (ja1):

I do but not fully

mathslover (mathslover):

Okay, so here your problem is rationalizing... We have : \(\cfrac{150r^2}{\sqrt{r^3}}\) We can write : \(\sqrt{r^3} \) as \((r^3)^\cfrac{1}{2} = r^{\cfrac{3}{2}} \) Right?

OpenStudy (ja1):

Woh wait what? LaTeX is all messed up for me :/

OpenStudy (luigi0210):

Now that Math puts it that way.. it is simple. Maybe sleep would do us some good huh JA?

OpenStudy (ja1):

Yes you have no idea....

mathslover (mathslover):